# Number Cookbook: Number Understanding of Language Models and How to Improve It

### v2

@@ A.1.3 Examples for each task
We provide each tasks with an examples. To test the models, we also add some model specific system messages like “You are a helpful assistant to process numbers. Please directly answer the question after the =”. The context before “$=$” is the question and the context after “$=$” is the groundtruth and be removed when testing.
@@
  • Multiply-Fraction: Multiply two numbers: $559.92e16\times 9.731e38=9.653152e$ 55
@@ A.1.4 Expected representation in each task
We summary the second number representation and result representation in each tasks in Table 4 and Table 5 where I means integer, i means (shorter) int, Fl means float, Fr means fraction, S means scientific notation and N means no such a number.
@@ A.1.5 Non-included tasks
    • Converting fractions to scientific notation can be done by first converting to a float. (Fraction-to_scientific = Fraction-to_float + Float-to_scientific). Fraction-SiginificantFigure is similar.
@@
    • count is a special task where we just consider a nubmer as “a set of digits” so count in a float, fraction and scientific notation is as the same as in a integer.
@@ A.1.7 Preprocess and question generation for NUPA tasks
After we select two random numbers, we have some pre-procession to generate the final questions:
@@
  • For “to_float-Fraction”, we require the fraction must can be convert into a finite decimal, that is the denominator contains only factors 2 and 5.
@@
The pre-procession could introduce additional duplicated data. we facilitate a post-filtering of duplication and decontainment
@@ A.4.1 Experiment details
Our experiments were conducted on a cluster equipped with Nvidia A800 GPUs (80GB memory). Training a 100M model takes 5–8 hours, a 1B model approximately 1 day, and a 3B model around 2 days on a single A800 GPU. Fine-tuning a pretrained model typically takes about 1 day.
@@ A.4.2 Tokenization
Figure 10: Accuracy of 0.9B models trained with 1-3 digit tokenizer on three task of integer addition, float addition and integer multiplication according to digit match. X-axis is the number of seen training samples.
@@
Figure 11: Accuracy of 0.9B models trained with 1-3 digit tokenizer on three task of integer addition, float addition and integer multiplication according to dlength. Here we report $log_{2}(\text{dlength}+1)$.X-axis is the number of seen training samples.
@@ A.4.3 PEs
We show exact match, digit match and dlength of 100M models trained with various PE, including RoPE, NoPE and Alibi in Figure 13. We find NoPE and Alibi achieve better length generalization than RoPE, which is consistent with previous work like Zhou et al. (2024).
@@ A.4.4 Data Formats
Previous work (Zhou et al., 2024) believes that reverse formatting can help the calculation of each digit by aligning the calculation order to the right-to-left order that humans are accustomed to and solve the carrying problem. That is, from left-to-right, we cannot determine the result of the current digit unless the next digit results and whether there is a carrying have been known. However, a more detailed analysis can explain why the order is not as important as previously believed:
@@
About the experiments of index hint, we show in Table 17. Our conclusion on index hints seems to contradict the findings of Zhou et al. (2024), where models with index hints appeared to achieve better results. We believe this discrepancy may be related to model size and digit range. In their work, a much smaller model (only 25M parameters) was used, but the training range covered 1-40 digits. This reduced the model’s ability to learn the patterns independently without external hints, resulting in a different learning outcome where the model began to rely on index hints. As a piece of evidence, when Zhou et al. (2024) train 1-10 digits, the performance without index hint is OK. (But they did not provide the complete results of 1-10 digit training in their work.) The effectiveness of index hints may involve complex interactions, which could be an interesting direction for future research.
@@
A.4.5 NUPA fintuning with PE, tokenizer and representation modification
@@ A.5.2 Maximum Digit Lengths within Context Window
The selective tasks used to train the RFFT are shown in Table 19 and we also report the maximal length within 2k tokens context windows limitation. For the detailed prompt of these tasks we cannot put them into papers so we include them in supplementary.

### v3

@@ A.1.3 Examples for each task
We provide each tasks with an example. To test the models, we also add some model specific system messages like “You are a helpful assistant to process numbers. Please directly answer the question after the =”. The context before “$=$” is the question and the context after “$=$” is the groundtruth and is removed when testing.
@@
  • Multiply-Scientific: Multiply two numbers: $559.92e16\times 9.731e38=9.653152e$ 55
@@ A.1.4 Expected representation in each task
We summarize the second number representation and result representation in each task in Table 4 and Table 5 where I means integer, i means (shorter) int, Fl means float, Fr means fraction, S means scientific notation and N means no such a number.
@@ A.1.5 Non-included tasks
    • Converting fractions to scientific notation can be done by first converting to a float. (Fraction-to_scientific = Fraction-to_float + Float-to_scientific). Fraction-SignificantFigure is similar.
@@
    • count is a special task where we just consider a number as “a set of digits” so count in a float, fraction and scientific notation is as the same as in a integer.
@@ A.1.7 Preprocess and question generation for NUPA tasks
After we select two random numbers, we have some preprocessing to generate the final questions:
@@
  • For “to_float-Fraction”, we require that the fraction can be converted into a finite decimal, that is the denominator contains only factors 2 and 5.
@@
The pre-processing could introduce additional duplicated data, so we implement a post-filtering step to remove duplicates and ensure data integrity.
@@ A.4.1 Experiment details
Our experiments were conducted on a cluster equipped with Nvidia A800 GPUs (80GB memory). Training a 100M model takes 5–8 hours, a 1B model approximately 1 day, and a 3B model around 2 days on a single A800 GPU. Finetuning a pretrained model typically takes about 1 day.
@@ A.4.2 Tokenization
Figure 10: Accuracy of 0.9B models trained with 1-3 digit tokenizer on three tasks of integer addition, float addition and integer multiplication according to digit match. X-axis is the number of seen training samples.
@@
Figure 11: Accuracy of 0.9B models trained with 1-3 digit tokenizer on three tasks of integer addition, float addition and integer multiplication according to dlength. Here we report $log_{2}(\text{dlength}+1)$.X-axis is the number of seen training samples.
@@ A.4.3 PEs
We show exact match, digit match and dlength of 100M models trained with various PE, including RoPE, NoPE and Alibi in Figure 13. We find NoPE and Alibi achieve better length generalization than RoPE, which is consistent with previous work like Zhou et al. (2024b).
@@ A.4.4 Data Formats
Previous work (Zhou et al., 2024b) believes that reverse formatting can help the calculation of each digit by aligning the calculation order to the right-to-left order that humans are accustomed to and solve the carrying problem. That is, from left-to-right, we cannot determine the result of the current digit unless the next digit results and whether there is a carrying have been known. However, a more detailed analysis can explain why the order is not as important as previously believed:
@@
About the experiments of index hint, we show in Table 17. Our conclusion on index hints seems to contradict the findings of Zhou et al. (2024b), where models with index hints appeared to achieve better results. We believe this discrepancy may be related to model size and digit range. In their work, a much smaller model (only 25M parameters) was used, but the training range covered 1-40 digits. This reduced the model’s ability to learn the patterns independently without external hints, resulting in a different learning outcome where the model began to rely on index hints. As a piece of evidence, when Zhou et al. (2024b) train 1-10 digits, the performance without index hint is OK. (But they did not provide the complete results of 1-10 digit training in their work.) The effectiveness of index hints may involve complex interactions, which could be an interesting direction for future research.
@@
A.4.5 NUPA finetuning with PE, tokenizer and representation modification
@@ A.5.2 Maximum Digit Lengths within Context Window
The selective tasks used to train the RFFT are shown in Table 19 and we also report the maximal length within 2k tokens context windows limitation.